As an Investment Banker your client, LMN Limited is looking at a restructuring its business. You extract the following data from the financials of the company.

Particulars Rs. Crs.

Equity capital (Face value Rs. 10) 1,000.00

Debentures (@ 12%) 400.00

Long Term unsecured loan (@15%) 200.00

Total 1,600.00


The company has been paying a dividend of 20% per annum (historically). The stock of LMN Limited is listed at Rs. 20 on the NSE. Compute the cost of:-

a. Equity capital.

b. Weighted Average cost of capital of LMN Limited


1
Expert's answer
2021-09-21T05:03:03-0400


a. Equity capital:

In accordance with the dividend approach to determining the cost of equity, the price of equity is determined by the current value of dividends that shareholders either receive or wish to receive from the company, as well as expected changes in the exchange value of shares. The logic of this approach is that the payment of dividends is considered as a payment for equity:

The cost of equity can be found as:

"the cost of Equity capital=\frac{interest rate on dividends\times100}{Equity capital}=\frac{20\times100}{10 000}=0.2"

Equity capital=0.2

b.

"WACC=keWe+kdWd"


"WACC=keWe+kdWd=0.2\times\frac{10000}{160000}+0.12\times\frac{4000}{160000}+0.15\times\frac{2000}{160000}=0.017" or 1.7%